OpenStudy (anonymous):

The formula for density D is given by D = m/v, where m is the mass of an object and v is its volume. Solve the formula for m. m =

OpenStudy (anonymous):

m.m?

OpenStudy (anonymous):

solve for m

OpenStudy (anonymous):

mutltiply bith sides by (v) D*v=v*(m/v) D*v=m since the (v) cancel each other